A rectangle has a perimeter of 42 feet. The length is 3 feet longer than the width . What is the lenght and width of the deck
Answer by sofiyac(983)

P=2(l+w)
if width is x then length is x+3
42=2(x+x+3)
42=2(2x+3)
21=2x+3
18=2x
9=x
so the width is 9 then the length is 9+3=12